OpenStudy (anonymous):

let h(x)=integral of 0 to x f(t)dt Find h(3), h'(3), and h"(3)

zepdrix (zepdrix):

Hmm this is a neat problem :) Fundamental Theorem of Calculus.

OpenStudy (anonymous):

\[h(x)=\int\limits_{0}^{x} f(t)dt \]

zepdrix (zepdrix):

We simply use FTC, Part 2 for the first part,\[h(3)=\int\limits\limits_{0}^{3} f(t)dt=F(t)|_0^3\]

zepdrix (zepdrix):

=F(3)-F(0) right? :)

zepdrix (zepdrix):

Were you given a function for f? :d

OpenStudy (kmeis002):

Graphically, what does \( \int_0^3 f(x) dx \) represent?